THE ROTATIONAL SPECTRUM OF TERTIARY-BUTYL ALCOHOL

\underline{E. A. COHEN}, B. J. DROUIN, Jet Propulsion Laboratory, California Institute of Technology, Pasadena, CA 91109-8099; E. A. VALENZUELA^{a}, R. C. WOODS, Department of Chemistry, University of Wisconsin, Madison WI 53706-1322; W. CAMINATI, A. MARIS, and S. MELANDRI, Dipartimento di Chimica “G. Ciamician” dell’Università, Via Selmi 2, I-40126 Bologna, Italy.

Tertiary-butyl alcohol is a nearly spherical rotor for which the internal rotation axis of the t-butyl group is close to the c molecular axis in the ac plane. Methyl group torsional spittings are not observed in the ground state. Its 8 to 40 GHz rotational spectrum was reported at this meeting by Valenzuela and Woods in 1974^{b} and in more detail in 1975^{c}. The parameters derived at that time from a fit to the E states with J, K ≤ 20^{d} have provided the basis for extending the measurements to > 500 GHz. The combined data set extends to J, K > 50 and is fit with the program spfit^{e} using a common set of parameters for both the A and E states. The general features of the spectrum and the fitting procedure will be described. The resulting molecular constants and their interpretation will be discussed.

aPresent address, 4901 Oakridge Dr. Midland, MI 48640-1919

bE.A. Valenzuela, and R. C. Woods, Abstract MF6, International Symposium on Molecular Spectroscopy, Columbus, OH, 1974

cE.A. Valenzuela, and R. C. Woods, Abstract RG15, International Symposium on Molecular Spectroscopy, Columbus, OH, 1975

dE.A. Valenzuela, Ph.D. Thesis, University of Wisconsin-Madison, 1975.

eH. M. Pickett, J. Mol. Spectrosc. 148 (1991) 271-377.